OpenStudy (anonymous):

solve using elimination 2x+3y=9 x+5y=8

OpenStudy (danjs):

multiply second equation by -2 add the 2 equations together no more x variable

OpenStudy (xapproachesinfinity):

multiply the second equation by -2 to get: -2x-10y=-16

OpenStudy (xapproachesinfinity):

you will have: 2x+3y=9 -2x-10y=-18

OpenStudy (xapproachesinfinity):

if you add those two equation like @DanJS said 2x+3y=9 -2x-10y=-18 =========== 0-7y=-9 solve for y once you have y look for x in one of the equations you started with